What would be the final concentration of a solution containing 20mL of 1.5% NaCl solution diluted to 200mL
1
Expert's answer
2020-10-05T14:31:26-0400


density of solution NaCl = 1.0089 g/ml

density of solution = m/V

m1 solution NaCl = density of solution NaCl * V1 solution NaCl = 1.0089 * 20 = 20.178 g

mass fraction = (m substance/ m solution) * 100%

m1 substance NaCl = mass fraction * m1 solution NaCl = 0.015 * 20.178 = 0.30267 g

V H2O = V2 solution - V1 solution = 200 - 20 = 180 ml

m H2O = density of H2O * V H2O = 1 * 180 = 180 g

mass fraction2 = m1 substance NaCl / (m1 solution NaCl + m H2O ) = 0.30267/(20.178 + 180) = 0.15%